OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = two thirdsx − 10 y = two thirdsx + 10 y = −three halvesx − 10 y = −three halvesx + 10

hartnn (hartnn):

can you find the slope of 2x-3y =3 ?

OpenStudy (haseeb96):

yes i can y=-3 and x = 2 ,c=-3 y=mx+c -3=m2-3 -3+3=2m 0\2=m m=0

OpenStudy (haseeb96):

hartnn i did

hartnn (hartnn):

thats incorrect...

OpenStudy (haseeb96):

how can i find it?

hartnn (hartnn):

isolate y from 2x − 3y = 3

OpenStudy (haseeb96):

-y=3-2x y=2x-3

OpenStudy (haseeb96):

-3y=3-2x y=-1+2\3 x here , b= -1 and m= +2\3 but we know that in perpendicular position we use m1m2=-1 m +2\3 =-1 m=-3\2 so m=-3\2 and the line passes from the (-8,2) (x,y)=(-8,2) y=mx+b 2=-8(-3\2)+b 2= 12+b b=2-12 b=-10 now put it in the slope intercept from y=-3\2x -10 this is the true answer it means that C option is absolutly correct

OpenStudy (shamil98):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). Perpendicular lines have negative reciprocal slopes. First get your equation in slope-intercept form (y=mx+b). by solving for y. 2x - 3y = 3 3y = 2x - 3 y = 2/3x - 1 the perpendicular slope is -3/2 Then plug in your (x,y) coordinates given and with your slope into. y = mx +b and Solve for the b value ( your y-intercept ) Do you understand so far?